A meter stick which is 100 cm is moving with an observed length of 60 cm. What is its speed?

1
Expert's answer
2022-05-02T13:42:04-0400
"l=l_0\sqrt{1-v^2/c^2}"

"v=c\sqrt{1-(l/l_0)^2}\\
=c\sqrt{1-(60/100)^2}=0.8c"